given a bar = i cap + 2 j cap +3 k cap and vector b = -i cap+ +2 j cap+ k cap and vector c = 3 i cap + j cap.  If resultant of these vectors and unit perpendicular to both a bar and b bar are given by 1/m(6 i cap + 10 j cap + 8 k cap ) and (- i cap - j cap + k cap)/n^(1/2) respectively , then m+n =
 

Grade:12th pass

1 Answers

Arun
5 years ago
Resultant vector = a + b + c = 3 i + 5j +4k
 
but if compare this to the given expression,
m = 2
 
Now
 
perpendicular vector to a and b = a X b = (2k – j + 2k + 2i – 3j – 6i) =  – 4i – 4j + 4k
 
unit vector =  – 4i – 4j + 4k / sqrt (48) = – i + – j + k / sqrt (3)
 
if we compare this to the given expression , then
 
n = 3
 
Hence
 
m + n = 2 +3 = 5